OpenStudy (anonymous):

Can someone please teach me how to solve this question.... Mr. Tobins taxable income is $8,900. On this, she must pay 2% on the first %3,000, 3% on teh next $3,000, and 4% on the balance for state income tax. During the year her employer had withheld $4.50 per week. To settle her tax bill at the end of the year, Mrs. Tobin had to pay an additional

OpenStudy (nubeer):

First find the Tax that has to be paid 2% of 3000 3% of 3000 4% of Remaining

OpenStudy (nubeer):

Can u do that?

OpenStudy (anonymous):

yes =60 =90 2900*0.04?=116

OpenStudy (anonymous):

@nubeer

OpenStudy (nubeer):

yes.. add all these up.. now her jobs take out 4.5$ per week.. so how much it would cost for an year?

OpenStudy (anonymous):

266-4.5?

OpenStudy (anonymous):

so if 4.50$ is withheld each week how would I subtract it for a year?

OpenStudy (anonymous):

@nubeer

OpenStudy (nubeer):

no.. 4.5 is with held per week. so how much they would with held in a year.. 4.5 x 4 = 18 ( 1 month amount) 18 x 12 = 216 with held in a year.. got it?

OpenStudy (anonymous):

yes, so 266-216=50

OpenStudy (nubeer):

yeah

OpenStudy (anonymous):

so.. mrs. tobin had to pay an additional 50$?

OpenStudy (anonymous):

That isn't the result according to my book, did I miss something?

OpenStudy (nubeer):

ok what's the answer according to your book..? i might have miss something

OpenStudy (anonymous):

I'm studying for a state math test

OpenStudy (anonymous):

its $32

OpenStudy (anonymous):

I've been stuck on this problem all day I tried this so many times and I keep getting different answers...

OpenStudy (anonymous):

Do you want to try helping me solve a similar problem and just ignore this one?

OpenStudy (nubeer):

ok sure..

OpenStudy (anonymous):

Thanks...

OpenStudy (anonymous):

Mr. Emerson sets aside $80 per month to cover realty and school taxes. His home is assessed at $2400. His realty tax is $3.14per $100, and his school tax is $1.19 per $100. The amount he must add at the end of the year to cover both taxes is..

OpenStudy (anonymous):

I multiplied 249*3.14=781.86 and 249*1.19 296.31

OpenStudy (anonymous):

I added them but I didn't get the correct result.. do I multiply 80*12 and subtract the results?0

OpenStudy (nubeer):

how you got 249?

OpenStudy (anonymous):

according the instructions in the book... since theres a hundred per 3.14 and 1.19 I should divide that by 24900

OpenStudy (nubeer):

in question you wrote 2400$ . :3

OpenStudy (anonymous):

oh my bad... typing mistake :3

OpenStudy (nubeer):

lol .. well yeah.. i think you are doing it right.. thats how it should be done.

OpenStudy (anonymous):

Why don't I get the correct result then??!! :(

OpenStudy (anonymous):

math hates me...I follow all the instructions and study for hours and I always get an incorrect answer...

OpenStudy (nubeer):

what is the correct answer in your book?

OpenStudy (nubeer):

yeah i know sometimes maths get pretty tough :3 but its really nice to see that you work really hard to learn. i am sure you will get the hang of it

OpenStudy (anonymous):

the correct answer is 313.02 I got 118.17 -_-

OpenStudy (anonymous):

Yeah, I hope so... and thanks.. I just hope I don't get a question like this on the test..

OpenStudy (anonymous):

well, thanks... for trying to help. Have a good night :)